# Diagonalization $of$ Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v31

@@ 1.1 Main Results
Thus, the computational complexity textbook [AB09] thinks of that these properties are also applicable for oracle Turing machines, and further regards that to show $\mathcal{P}\neq\mathcal{NP}$ requires some other properties in addition to the properties I and II as stated above. However, we would like to say that this kind of perspective is not absolutely (fully) correct either. Notice also that similar point of view was followed in [Sip13, BC94], i.e., the references [Sip13] and [BC94] hold the viewpoint that if we could prove that $\mathcal{P}$ and $\mathcal{NP}$ were different by diagonalizing, then we could conclude that they are different relative to any oracle as well. Interestingly enough, we will demonstrate the key points besides the aforementioned assumptions I and II when diagonalization technique is applicable such that we can overcome the so-called “Relativization Barrier.”
But at present, for convenience of the reader, let us first quote the notion of an enumeration of a set as follows:
Definition 1.1 ([Rud76], p. 27, Definition 2.7)
555 In Cantor’s terminology, the enumeration of something is the “sequence” of something. We should be clear that only enumerable sets have enumerations. And by the term “enumerable”, Turing refers to [Hob21], p. 78. That is, an aggregate (i.e., set) which contains an indefinitely great number of elements is said to be enumerable, or countable, when the aggregate is such that a $(1,1)$ correspondence can be established between the elements and the set of integral numbers $1,2,3,\cdots,$ i.e., $\mathbb{N}_{1}$. We can simply deem that an enumeration of an enumerable set $T$ is just a function $e:\mathbb{N}_{1}\rightarrow T$ which is surjective; or equivalently, is an injective function $e^{\prime}:T\rightarrow\mathbb{N}_{1},$ meaning that every element in $T$ corresponds to a different element in $\mathbb{N}_{1}$. See [Tur37], Section of Enumeration of computable sequences.
@@
To convince the experts who argue that the “Relativization Barrier” is a real barrier that should overcome when proving $\mathcal{P}\neq\mathcal{NP}$, we show the following important theorem that concerns oracle Turing machines, which is on purpose to demonstrate that the “Relativization Barrier” is not really a barrier. To do so, of course, we should first suppose without loss of generality that polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings (i.e., the above property I), and further, there are universal nondeterministic oracle Turing machines that can simulate and flip the answers of other deterministic oracle Turing machines without much overhead in running time or space (i.e., the above property II). Indeed, the next theorem being introduced in the following is our conclusions in this regard.
@@
Theorem 5
666 The proof of this theorem, in fact, is similar to the proof of Cantor’s theorem: there are infinite sets that can not be put into one-to-one correspondence with the set of positive integers, i.e., $\mathbb{N}_{1}$; see e.g. [Gra94]. Furthermore, the argument of this theorem lies in the assumptions that: (I) polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings; (II) a universal nondeterministic oracle Turing machine exists that can simulate and flip the answers of other deterministic oracle Turing machines; and (III) the simulation of a universal nondeterministic oracle Turing machine to any deterministic oracle Turing machine can be done within $O(T(n)\log T(n))$ steps, where $T(n)$ is the time complexity of the simulated deterministic oracle Turing machine.
@@
If a set $T$ (of oracle Turing machines) is enumerable, then the diagonalization technique may be applicable. In other words, that $T$ (of oracle Turing machines) is enumerable is an important prerequisite for the application of diagonalization technique.
1.2 Our Approach
@@
The Cook-Levin Theorem is a well-known theorem stating that Satisfiability (SAT) is complete for the complexity class $\mathcal{NP}$ under polynomial-time many-one reductions. Starting from this, on the one hand, if one wants to prove that $\mathcal{P}$ and $\mathcal{NP}$ are identical, then she/he can try to design polynomial-time algorithms for SAT. On the other hand, if one wants to show that $\mathcal{P}$ and $\mathcal{NP}$ differ, then she/he may try to prove a super-polynomial lower-bound for Satisfiability.
@@
As a novel idea and attempt, we enumerate all polynomial-time deterministic Turing machines and then diagonalize against all of them with a universal nondeterministic Turing machine, thus obtaining a language $L_{d}$ not in $\mathcal{P}$. Based on the above work, we then carefully do an analysis for the language $L_{d}$ showing that this language is in $\mathcal{NP}$. To stress that it is “novel” is because, in general, a common practice is to use a universal deterministic Turing machine to diagonalize against a list of deterministic Turing machines (see e.g. [Pap94, AB09]), or to employ a universal nondeterministic Turing machine to diagonalize against a list of nondeterministic Turing machines; see e.g. [AB09, For00, FS07, Zak83]. That using a universal nondeterministic Turing machine to diagonalize against all of the polynomial-time deterministic Turing machines is less common and can be see as a new attempt. We would like to emphasize that this approach has seldom been tried before in the literature, except that recently we noted that in [Sip92], Sipser mentioned an idea (see [Sip92], p. 605), this argument of which fails, to give a nondeterministic polynomial-time Turing machine which has an opportunity to run each of the deterministic polynomial-time Turing machines and arrange to accept a differing language.
@@
We should also point out that our idea of diagonalization against all of the polynomial-time deterministic Turing machines by using a universal nondeterministic Turing machine was inspired by our recent work [Lin21a]. Objectively, the work [Lin21a] is the source of our understanding in-depth and application of diagonalization technique in the domain of computational complexity, due to that the idea of diagonalization against deterministic Turing machines with a nondeterministic Turing machine goes back to [Lin21a], in which we use a universal nondeterministic $n$ space-bounded Turing machine to diagonalize against a collection of deterministic $n$ space-bounded Turing machines.
@@
As we mentioned earlier, we do not know whether the circuit lower bounds [Raz85] has really a natural proofs barrier to the $\mathcal{P}$ versus $\mathcal{NP}$ problem. Anyway, even that some day there were a proof of that $\mathcal{NP}\not\subseteq\mathcal{P}/poly$ by circuit lower bounds, we still believe that the approach presented in this work is the simplest, since we do not like complicated proofs either or we do not like to make things much more complicated.
1.3 Related Work
@@
As is well known, a central open question in computational complexity theory is the $\mathcal{P}$ versus $\mathcal{NP}$ problem, which is to determine whether every language accepted by some nondeterministic Turing machine in polynomial-time is also accepted by some deterministic Turing machine in polynomial-time. In this subsection, we will review its history and related works. With regard to the importance of the problem, we refer the reader to the references [Coo00, Coo03, Wig07].
@@
A year later, stimulated by the work of Cook [Coo71] (according to the viewpoints in [Kar72]), Karp [Kar72] used these completeness results to show the celebrated conclusions that $20$ other natural problems are $\mathcal{NP}$-complete, forcefully demonstrating the importance of the subject. Thus far, there are many problems shown to be $\mathcal{NP}$-complete. See the excellent reference [GJ79] on this subject. In his paper [Kar72], Karp also introduced the now standard notation $\mathcal{P}$ and $\mathcal{NP}$ and redefined $\mathcal{NP}$-completeness by using the polynomial-time analog of many-one reducibility, which has become standard. Meanwhile, Levin [Lev73], independently of Cook [Coo71] and Karp [Kar72], defined the notion of “universal search problem”, similar to the $\mathcal{NP}$-complete problem, and gave six examples, which include Satisfiability.
Although the precise statement of $\mathcal{P}$ versus $\mathcal{NP}$ question was formally defined in the $1970$s in his seminal paper [Coo71] by Cook, there were previous inklings of the problems involved (see [A1]). A mention of the underlying problem occurred in a $1956$ letter written by K. Gödel to J. von Neumann. Gödel asked whether theorem-proving could be solved in quadratic or linear time (see e.g. [Har89]). It is worth paying attention that, besides the classical version of the question, there is a version expressed in terms of the field of complex numbers, which has attracted the interest of the mathematics community [BCSS98].
@@
1.4 Overview
The rest of the work is organized as follows: For the convenience of the reader, in the next Section we will review some notions closely associated with our discussions and fix some notation we will use in the following context. Also, some useful technical lemmas are presented. In Section 3, we provide a method to encode a polynomial-time deterministic Turing machine to an integer so that we can prove that the set of all polynomial-time deterministic Turing machines is enumerable. Section 4 contains the construction of our nondeterministic Turing machine which accepts a language $L_{d}$ not in $\mathcal{P}$. And Section 5 overcomes the obstacle to prove the language accepted by our simulating machine is in $\mathcal{NP}$. Section 6 is devoted to prove Theorem 4, which states that there is a language $L_{s}\in\mathcal{P}$ and the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. In Section 7, we focus our attention on overcoming the so-called “Relativization Barrier.” Finally, we draw some conclusions in the last Section.
2 Preliminaries
In this section, we describe the notation and notions needed in the following context. We would like to point out that the our style of writing888 But merely the style of writing. Indeed, if to describe the preliminaries in author’s own words it will appear rather verbose. in this Section is heavily influenced by that in in Aho, Hopcroft and Ullman’s book [AHU74].
@@ 2 Preliminaries
i.e., the positive integers. It is clear that there is a bijection between $\mathbb{N}$ and $\mathbb{N}_{1}$. To see this just let the bijection to be
$n\mapsto n+1$
@@
The computation model we use here is the Turing machine as it defined in standard textbooks such as [HU69, HMU06]. Here, we follow the standard definition presented in [AHU74]:
Definition 2.1 ($k$-tape deterministic Turing machine, p. $26$, $27$ in [AHU74])
@@
  7. 7.
    $\delta$ is the next-move function, maps a subset of $Q\times T^{k}$ to
@@
Definition 2.2 ($k$-tape nondeterministic Turing machine, p. $365$ in [AHU74])
@@
Let $M(w)$ denote that Turing machine $M$ is on input $w$. If for every input $w$ of length $n$,999 In the following context, we will use $|w|$ to denote the length of $w$. all computations of $M$ end in less than or equal to $T(n)$ steps, then $M$ is said to be a deterministic (respectively, nondeterministic) $T(n)$ time-bounded Turing machine, or is said to be of time complexity $T(n)$.
The family of languages of deterministic time complexity $T(n)$ is denoted by DTIME[$T(n)$]; the family of languages of nondeterministic time complexity $T(n)$ is denoted by NTIME[$T(n)$]. The notation $\mathcal{P}$ and $\mathcal{NP}$ are defined to be the class of languages:
@@
The above definitions on complexity classes $\mathcal{P}$ and $\mathcal{NP}$ look a bit different from the official standard definitions in [Coo00]. However, we will show in Appendix A that they are in fact equivalent.
With respect to the time complexity of $k$-tape nondeterministic (respectively, deterministic) Turing machine and that of single-tape nondeterministic (respectively, deterministic) Turing machine, we have the following useful lemma and corollary, extracted from [AHU74] (see Lemma 10.1 and Corollary 1 to Lemma 10.1 in [AHU74]), which play important roles in the following context:
Lemma 2.1 (Lemma 10.1 in [AHU74])
101010 The deterministic version of this theorem to appear in [HS65] for the first time (Theorem 6 in [HS65]), and one can extend its proof to show the nondeterministic version of this lemma.
@@
Corollary 7 (Corollary 1 in [AHU74] to Lemma 10.1; see also Theorem 6 in [HS65] and Theorem 2.1 in [Pap94])
111111 This corollary immediately follows from the Lemma 10, since the $k$-tape deterministic Turing machine is a special kind of $k$-tape nondeterministic Turing machine.
@@
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$. For example, ${\rm co}\mathcal{P}$ is the complement of $\mathcal{P}$, and ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that, the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa.
@@
3 Enumeration of All Polynomial-Time Deterministic Turing Machines
We should first clearly clarify the formal definition about polynomial-time deterministic Turing machines before coming to the point.
Definition 3.1 (Cf. [Coo00])
@@ 3 Enumeration of All Polynomial-Time Deterministic Turing Machines
Remark 3.1
Obviously, in the above definition, given a polynomial-time deterministic Turing machine $(M,k)$, for any input $x$ of length $|x|$, $M(x)$ will halt within $O(|x|^{k+i})$ steps, where $i\geq 0$. But there exists some input $y$ of length $|y|$, $M(y)$ does not halt within $O(|y|^{k-1})$ steps. Further, the tuple-representation $(M,k)$ of a deterministic $n^{k}+k$ time-bounded Turing machine $M$, due to this work, has some advantages. Namely, it clearly indicates that $M$ is a polynomial-time deterministic Turing machine and that the order of $(M,k)$ is $k$, from which we can easily recover the minimal polynomial $n^{k}+k$.
By Corollary 11, we can restrict ourselves to single-tape deterministic Turing machines. So, in the following context, by polynomial-time deterministic Turing machines we mean single-tape polynomial-time deterministic Turing machines.
To obtain our main result, we need to enumerate the polynomial-time deterministic Turing machines, so that for each nonnegative integer $i$ there is a unique tuple of $(M,k)$ associated with $i$ (i.e., to define a function from $\mathbb{N}_{1}$ to the set of all polynomial-time deterministic Turing machines $\{(M,k)\}$ such that the function is surjective131313 There are a variety of ways to enumerate all polynomial-time deterministic Turing machines. For instance, see proof of Theorem 14.1 in [Pap94], p. 330; or see [Lad75].), such that we can refer to the $j$-th polynomial-time deterministic Turing machine.
To achieve our goals, we first use the method presented in [AHU74], p. 407, to encode a single-tape deterministic Turing machine into an integer.141414 The way to enumerate deterministic Turing machines in [AHU74] is basically the same as that of Turing [Tur37], i.e., to define a mapping from Turing machines to positive integers, but with different specific details. See [Tur37], Section of Enumeration of computable sequences.
@@
Any integer that cannot be decoded is deemed to represent the trivial polynomial-time deterministic Turing machine with an empty next-move function by this encoding. Every polynomial-time deterministic Turing machine will appear infinitely often in the enumeration since, given a polynomial-time deterministic Turing machine, we can prefix $1$’s at will to find larger and larger integers representing the same set of the polynomial-time deterministic Turing machine of $(M,k)$. We denote such a set of the polynomial-time deterministic Turing machine by $\widehat{M}_{j}$, where $j$ is the integer representing $(M,k)$. The reader will easily get that we have defined a surjective function $e$ from $\mathbb{N}_{1}$ to the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines, which is consistent with the Definition 5.
Furthermore, we in fact have defined a $(1,1)$ correspondence between the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines and $\mathbb{N}_{1}$ if any integer that cannot be decoded is deemed to represent the trivial polynomial-time deterministic Turing machine, from which we have reached the similar case to p. 241 of [Tur37], i.e., the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines are therefore enumerable.
Remark 3.2
One of the conveniences of tuple-representation $(M,k)$ of a polynomial-time deterministic Turing machine in this way is, of course, to conveniently control the running time of the universal nondeterministic Turing machine $M_{0}$ constructed in Theorem 9 in Section 4 below, so that it facilitates our analysis of the time complexity of $M_{0}$, i.e., to easily show the fact of Theorem 10.
@@
Theorem 8
151515 This theorem is somewhat redundant because an enumeration of the set $\{(M,k)\}$ is a surjective function $e:\mathbb{N}_{1}\rightarrow\{(M,k)\},$ so that for any element $(M,k)$, there is an $i\in\mathbb{N}_{1}$ such that $e(i)=(M,k).$ For simplicity, we just let $P$ denote the set $\{(M,k)\}$ of polynomial-time deterministic Turing machines.
All of the polynomial-time deterministic Turing machines are in the above enumeration $e$. In other words, the set $P$ of all polynomial-time deterministic Turing machines is enumerable.
Remark 3.3

### v33

@@ 1.1. Main Results
Thus, the computational complexity textbook [AB09] thinks of these properties as also applicable for oracle Turing machines and further regards that to show $\mathcal{P}\neq\mathcal{NP}$ requires some other properties in addition to the properties I and II as stated above. However, we would like to say that this kind of perspective is not absolutely (fully) correct either. Notice also that a similar point of view was followed in [Sip13, BC94], i.e., the references [Sip13] and [BC94] hold the viewpoint that if we could prove that $\mathcal{P}$ and $\mathcal{NP}$ were different by diagonalizing, then we could conclude that they are different relative to any oracle as well. Interestingly enough, we will demonstrate the key points besides the aforementioned assumptions I and II when diagonalization techniques are applicable such that we can overcome the so-called “Relativization Barrier.”
But at present, for the convenience of the reader, let us first quote the notion of an enumeration of a set as follows:
Definition 1.5 ([Rud76], p. 27, Definition 2.7).
555 In Cantor’s terminology, the enumeration of something is the “sequence” of something. We should be clear that only enumerable sets have enumerations. And by the term “enumerable”, Turing refers to [Hob21], p. 78. That is, an aggregate (i.e., set) that contains an indefinitely great number of elements is said to be enumerable, or countable, when the aggregate is such that a $(1,1)$ correspondence can be established between the elements and the set of integral numbers $1,2,3,\cdots,$ i.e., $\mathbb{N}_{1}$. We can simply deem that an enumeration of an enumerable set $T$ is just a function $e:\mathbb{N}_{1}\rightarrow T$ that is surjective; or equivalently, it is an injective function $e^{\prime}:T\rightarrow\mathbb{N}_{1},$ meaning that every element in $T$ corresponds to a different element in $\mathbb{N}_{1}$. See [Tur37], Section of Enumeration of computable sequences.
@@
To convince the experts who argue that the “Relativization Barrier” is a real barrier that should be overcome when proving $\mathcal{P}\neq\mathcal{NP}$, we show the following important theorem that concerns oracle Turing machines, which is on purpose to demonstrate that the “Relativization Barrier” is not really a barrier. To do so, of course, we should first suppose without loss of generality that polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings (i.e., the above property I), and further, there are universal nondeterministic oracle Turing machines that can simulate and flip the answers of other deterministic oracle Turing machines without much overhead in running time or space (i.e., the above property II). Indeed, the next theorem being introduced in the following is our conclusion in this regard.
@@
Theorem 1.6.
666 The proof of this theorem, in fact, is similar to the proof of Cantor’s theorem: there are infinite sets that can not be put into one-to-one correspondence with the set of positive integers, i.e., $\mathbb{N}_{1}$; see e.g., [Gra94]. Furthermore, the argument of this theorem lies in the assumptions that (I) polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings; (II) a universal nondeterministic oracle Turing machine exists that can simulate and flip the answers of other deterministic oracle Turing machines; and (III) the simulation of a universal nondeterministic oracle Turing machine to any deterministic oracle Turing machine can be done within $O(T(n)\log T(n))$ steps, where $T(n)$ is the time complexity of the simulated deterministic oracle Turing machine.
@@
If a set $T$ (of oracle Turing machines) is enumerable, then the diagonalization technique may be applicable. In other words, that $T$ (of oracle Turing machines) is enumerable is an important prerequisite for the application of diagonalization techniques.
1.2. Our Approach
@@
The Cook-Levin Theorem is a well-known theorem stating that Satisfiability (SAT) is complete for the complexity class $\mathcal{NP}$ under polynomial-time many-one reductions. Starting from this, on the one hand, if one wants to prove that $\mathcal{P}$ and $\mathcal{NP}$ are identical, then she/he can try to design polynomial-time algorithms for SAT. On the other hand, if one wants to show that $\mathcal{P}$ and $\mathcal{NP}$ differ, then she/he may try to prove a super-polynomial lower bound for Satisfiability.
@@
As a novel idea and attempt, we enumerate all polynomial-time deterministic Turing machines and then diagonalize against all of them with a universal nondeterministic Turing machine, thus obtaining a language $L_{d}$ not in $\mathcal{P}$. Based on the above work, we then carefully do an analysis for the language $L_{d}$, showing that this language is in $\mathcal{NP}$. To stress that it is “novel” is because, in general, a common practice is to use a universal deterministic Turing machine to diagonalize against a list of deterministic Turing machines (see e.g., [Pap94, AB09]), or to employ a universal nondeterministic Turing machine to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, Zak83]. That using a universal nondeterministic Turing machine to diagonalize against all of the polynomial-time deterministic Turing machines is less common and can be seen as a new attempt. We would like to emphasize that this approach has seldom been tried before in the literature, except that recently we noted that in [Sip92], Sipser mentioned an idea (but this simple argument fails; see [Sip92], p. 605) to give a nondeterministic polynomial-time Turing machine that has an opportunity to run each of the deterministic polynomial-time Turing machines and arrange to accept a differing language.
@@
We should also point out that our idea of diagonalization against all of the polynomial-time deterministic Turing machines by using a universal nondeterministic Turing machine was inspired by our recent work [Lin21]. Objectively, the work [Lin21] is the source of our in-depth understanding and application of diagonalization techniques in the domain of computational complexity, due to the fact that the idea of diagonalization against deterministic Turing machines with a nondeterministic Turing machine goes back to [Lin21], in which we use a universal nondeterministic $n$ space-bounded Turing machine to diagonalize against a collection of deterministic $n$ space-bounded Turing machines.
@@
As we mentioned earlier, we do not know whether the circuit lower bounds [Raz85] really have a natural proofs barrier to the $\mathcal{P}$ versus $\mathcal{NP}$ problem. Anyway, even if someday there were proof of that $\mathcal{NP}\not\subseteq\mathcal{P}/poly$ by circuit lower bounds, we still believe that the approach presented in this work is the simplest, since we do not like complicated proofs either, nor do we like to make things much more complicated.
1.3. Related Work
@@
As is well known, a central open question in computational complexity theory is the $\mathcal{P}$ versus $\mathcal{NP}$ problem, which is to determine whether every language accepted by some nondeterministic Turing machine in polynomial time is also accepted by some deterministic Turing machine in polynomial time. In this subsection, we will review its history and related works. With regard to the importance of the problem, we refer the reader to the references [Coo00, Coo03, Wig07].
@@
A year later, stimulated by the work of Cook [Coo71] (according to the viewpoints in [Kar72]), Karp [Kar72] used these completeness results to show the celebrated conclusions that $20$ other natural problems are $\mathcal{NP}$-complete, forcefully demonstrating the importance of the subject. Thus far, there are many problems shown to be $\mathcal{NP}$-complete. See the excellent reference [GJ79] on this subject. In his paper [Kar72], Karp also introduced the now standard notation $\mathcal{P}$ and $\mathcal{NP}$ and redefined $\mathcal{NP}$-completeness by using the polynomial-time analog of many-one reducibility, which has become standard. Meanwhile, Levin [Lev73], independently of Cook [Coo71] and Karp [Kar72], defined the notion of “universal search problem,” similar to the $\mathcal{NP}$-complete problem, and gave six examples, which include Satisfiability.
Although the precise statement of the $\mathcal{P}$ versus $\mathcal{NP}$ question was formally defined in the $1970$s in his seminal paper [Coo71] by Cook, there were previous inklings of the problems involved (see [A1]). A mention of the underlying problem occurred in a $1956$ letter written by K. Gödel to J. von Neumann. Gödel asked whether theorem-proving could be solved in quadratic or linear time (see e.g., [Har89]). It is worth paying attention that, besides the classical version of the question, there is a version expressed in terms of the field of complex numbers, which has attracted the interest of the mathematics community [BCSS98].
@@
1.4. Organization
The rest of the work is organized as follows: For the convenience of the reader, in the next section we will review some notions closely associated with our discussions and fix some notation we will use in the following context. Also, some useful technical lemmas are presented. In Section 3, we provide a method to encode a polynomial-time deterministic Turing machine to an integer so that we can prove that the set of all polynomial-time deterministic Turing machines is enumerable. Section 4 contains the construction of our nondeterministic Turing machine, which accepts a language $L_{d}$ not in $\mathcal{P}$. And Section 5 overcomes the obstacle to prove the language accepted by our simulating machine is in $\mathcal{NP}$. Section 6 is devoted to proving Theorem 1.4, which states that there is a language $L_{s}\in\mathcal{P}$ and the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. In Section 7, we focus our attention on overcoming the so-called “Relativization Barrier.” Finally, we draw some conclusions in the last section.
2. Preliminaries
In this section, we describe the notation and notions needed in the following context. We would like to point out that our style of writing888 But merely the style of writing. Indeed, if we describe the preliminaries in the author’s own words, it will appear rather verbose. in this section is heavily influenced by that in Aho, Hopcroft, and Ullman’s book [AHU74].
@@ 2. Preliminaries
i.e., the positive integers. It is clear that there is a bijection between $\mathbb{N}$ and $\mathbb{N}_{1}$. To see this, just let the bijection be
$n\mapsto n+1,$
@@
The computation model we use here is the Turing machines as it is defined in standard textbooks such as [HU69, HMU06]. Here, we follow the standard definition presented in [AHU74]:
Definition 2.1 ($k$-tape deterministic Turing machine, p. $26$, $27$ in [AHU74]).
@@
  7. (7)
    $\delta$ is the next-move function, which maps a subset of $Q\times T^{k}$ to
@@
Definition 2.2 ($k$-tape nondeterministic Turing machine, p. $365$ in [AHU74]).
@@
Let $M(w)$ denote that Turing machine $M$ is on input $w$. If for every input $w$ of length $n$, [^9] In the following context, we will use $|w|$ to denote the length of $w$. all computations of $M$ end in less than or equal to $T(n)$ steps, then $M$ is said to be a deterministic (respectively, nondeterministic) $T(n)$ time-bounded Turing machine, or is said to be of time complexity $T(n)$.
The family of languages of deterministic time complexity $T(n)$ is denoted by ${\rm DTIME}[T(n)]$; the family of languages of nondeterministic time complexity $T(n)$ is denoted by ${\rm NTIME}[T(n)]$. The notation $\mathcal{P}$ and $\mathcal{NP}$ is defined respectively to be the class of languages:
@@
The above definitions of complexity classes $\mathcal{P}$ and $\mathcal{NP}$ look a bit different from the official standard definitions in [Coo00]. However, we will show in Appendix LABEL:sec:appendix that they are in fact equivalent.
With respect to the time complexity of a $k$-tape nondeterministic (respectively, deterministic) Turing machine and that of a single-tape nondeterministic (respectively, deterministic) Turing machine, we have the following useful lemma and corollary, extracted from [AHU74] (see Lemma 10.1 and Corollary 1 to Lemma 10.1 in [AHU74]), which play important roles in the following context:
Lemma 2.3 (Lemma 10.1 in [AHU74]).
101010 The deterministic version of this theorem appeared in [HS65] for the first time (Theorem 6 in [HS65]), and one can extend its proof to show the nondeterministic version of this lemma.
@@
Corollary 2.4 (Corollary 1 in [AHU74] to Lemma 10.1; see also Theorem 6 in [HS65] and Theorem 2.1 in [Pap94]).
111111 This corollary immediately follows from Lemma 10, since a $k$-tape deterministic Turing machine is a special kind of $k$-tape nondeterministic Turing machine.
@@
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$. For example, ${\rm co}\mathcal{P}$ is the complement of $\mathcal{P}$, and ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa.
@@ 3. Enumeration of All Polynomial-Time Deterministic Turing Machines
We should first clearly clarify the formal definition of polynomial-time deterministic Turing machines before coming to the point.
Definition 3.1 (Cf. [Coo00]).
@@
Remark 3.2.
Obviously, in the above definition, given a polynomial-time deterministic Turing machine $(M,k)$, for any input $x$ of length $|x|$, $M(x)$ will halt within $O(|x|^{k+i})$ steps, where $i\geq 0$. But there exists some input $y$ of length $|y|$ such that $M(y)$ does not halt within $O(|y|^{k-1})$ steps. Further, the tuple representation $(M,k)$ of a deterministic $n^{k}+k$ time-bounded Turing machine $M$ has some advantages in this work. Namely, it clearly indicates that $M$ is a polynomial-time deterministic Turing machine and that the order of $(M,k)$ is $k$, from which we can easily recover the minimal polynomial $n^{k}+k$.
By Corollary 11, we can restrict ourselves to single-tape deterministic Turing machines. So, in the following context, by polynomial-time deterministic Turing machines, we mean single-tape polynomial-time deterministic Turing machines.
To obtain our main results, we need to enumerate the polynomial-time deterministic Turing machines so that for each positive integer $i$ there is a unique tuple of $(M,k)$ associated with $i$ (i.e., to define a function from $\mathbb{N}_{1}$ to the set of all polynomial-time deterministic Turing machines $\{(M,k)\}$ such that the function is surjective 131313 There are a variety of ways to enumerate all polynomial-time deterministic Turing machines. For instance, see proof of Theorem 14.1 in [Pap94], p. 330, or see [Lad75].) such that we can refer to the $j$-th polynomial-time deterministic Turing machine.
To achieve our goals, we first use the method presented in [AHU74], p. 407, to encode a single-tape deterministic Turing machine into an integer.141414 The way to enumerate deterministic Turing machines in [AHU74] is basically the same as that of Turing [Tur37] in principle, i.e., to define a mapping from Turing machines to positive integers, but with different specific details. See [Tur37], Section of Enumeration of computable sequences.
@@
Any integer that cannot be decoded is deemed to represent the trivial polynomial-time deterministic Turing machine with an empty next-move function by this encoding. Every polynomial-time deterministic Turing machine will appear infinitely often in the enumeration since, given a polynomial-time deterministic Turing machine, we can prefix $1$’s at will to find larger and larger integers representing the same set of the polynomial-time deterministic Turing machine of $(M,k)$. We denote such a set of the polynomial-time deterministic Turing machine by $\widehat{M}_{j}$, where $j$ is the integer representing $(M,k)$. The reader will easily get that we have defined a surjective function $e$ from $\mathbb{N}_{1}$ to the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines, which is consistent with Definition 5.
Furthermore, we in fact have defined a $(1,1)$ correspondence between the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines and $\mathbb{N}_{1}$ if any integer that cannot be decoded is deemed to represent the trivial polynomial-time deterministic Turing machine, from which we have reached the similar case to p. 241 of [Tur37], i.e., the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines is therefore enumerable.
Remark 3.3.
One of the conveniences of tuple representation $(M,k)$ for a polynomial-time deterministic Turing machine in this way is, of course, to conveniently control the running time of the universal nondeterministic Turing machine $M_{0}$ constructed in Theorem 4.1 in Section 4 below, so that it facilitates our analysis of the time complexity of $M_{0}$, i.e., to easily show the fact of Theorem 4.4.
@@
Theorem 3.4.
151515 This theorem is somewhat redundant because an enumeration of the set $\{(M,k)\}$ is a surjective function $e:\mathbb{N}_{1}\rightarrow\{(M,k)\},$ so that for any element $(M,k)$, there is an $i\in\mathbb{N}_{1}$ such that $e(i)=(M,k).$ For simplicity, sometimes we just let $P$ denote the set $\{(M,k)\}$ of polynomial-time deterministic Turing machines.
All of the polynomial-time deterministic Turing machines are in the above enumeration $e$. In other words, the set $\{(M,k)\}$ of all polynomial-time deterministic Turing machines is enumerable.
Remark 3.5.